OpenStudy (anonymous):

PLS HELP!!!!!!!!!!! Solve the radical inequality. square root (x + 2) less than equal to x- 10

OpenStudy (kainui):

Try to square both sides. Also, ignore solutions (if any) that do not satisfy the original equation.

OpenStudy (anonymous):



OpenStudy (anonymous):

how do i solve frm there

OpenStudy (kainui):

You should write the inequality like this first.\[0 \le \sqrt{x-2} \le x - 10\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

then do i square both sides?

OpenStudy (kainui):

Yes, you do.

OpenStudy (anonymous):

can u pls share what you get when you square both sides

OpenStudy (kainui):

\[(\sqrt{x+2})^2 \le (x -10)^2\]\[x + 2 \le x^2 - 20x + 100\]

OpenStudy (anonymous):